# ../Figures/Fridge/

### v1

@@ Abstract
Magnetic impurities are known to degrade superconductivity. For this reason, physical vapor deposition chambers that have previously been used for magnetic materials have generally been avoided for making high-quality superconducting resonator devices. In this article, we show by example that such chambers can be used: with Nb films sputtered in a chamber that continues to be used for magnetic materials, we demonstrate compact ( gap) coplanar waveguide resonators with low-power internal quality factors near one million. We achieve this using a resist strip bath with no post-fabrication acid treatment, which results in performance comparable to previous strip baths with acid treatments. We also find evidence that this improved resist strip bath provides a better surface chemical template for post-fabrication hydrogen fluoride processing. These results are consistent across three Si substrate preparation methods, including a anneal.
@@ I Introduction
Superconducting microwave circuits are one of the leading candidate platforms for enabling quantum computing technology [kjaergaard_superconducting_2020, sivak_real-time_2023, acharya_quantum_2024], and Nb on Si is a commonly implemented materials set [mcrae_materials_2020]. A main bottleneck for improving device performance and technological prospects is to eliminate defects from the materials and interfaces within the superconducting circuits [mcrae_materials_2020, de_leon_materials_2021].
@@ III.1 Si preparation series
We use the open-source resonator design developed by the Boulder Cryogenic Quantum Testbed to characterize the microwave properties of resonators fabricated from these films [kopas_simple_2022]. This design has a gap between the center conductor and the ground plane and has a hangar resonator configuration, with eight resonators at frequencies between 4 and . For each resonance, we locally vary the probe frequency and fit the complex-valued transfer function to the expected resonance response, as in Figure 2(a). Accomplishing this as a function of drive power, and using the measured attenuation of the input coaxial lines, we can estimate the average photon number. We fit the value of the internal quality factor (details in Appendix D.2) and calculate the loss ():
@@
In Figure 2(b), we show the internal loss of a resonator as a function of the average photon number in the resonator for devices corresponding to our three substrate preparation methods. The three sample resonators plotted are representative of the qualitatively different trends observed. In particular, the data from resonator 2 and resonator 3 show signs of a ‘double-S’ shaped curve, possibly indicative of two TLS (two level system) baths existing at different critical photon numbers [verjauw_investigation_2021]. This suggests that a usual TLS-based model is unsuitable for our data. Note that the data from resonator 1 does not have a second plateau present in the 1 to 100 photon range. Furthermore, none of the resonators loss curves saturate at high powers. One possible explanation is the renormalization of the critical exponent for the TLS bath due to TLS-TLS interactions, but this does not completely avoid a saturation effect [crowley_disentangling_2023, altoe_localization_2022, verjauw_investigation_2021]. Another possibility is a large number of TLS baths with a wide distribution of critical photon numbers. Given these discrepancies with usual models, we do not attempt to fit to those models.
@@
The results of the substrate preparation series are summarized in Figure 2(c). We believe that the distributions across all the dies from the different substrate preparation methods are not sufficiently well-separated to draw a clear distinction. As Nb resonators are susceptible to significant losses from the reformation of Nb oxide [zheng_nitrogen_2022, verjauw_investigation_2021, altoe_localization_2022, woods_determining_2019], some of the variation in the data may be explained by the varying amount of Nb re-oxidation after the HF treatment. In addition, temporal fluctuations of measured quality factor during the same cooldown for Nb on Si were found to have a relative standard deviation of 13% [vallieres_loss_2024], much smaller than what we observe cooldown-to-cooldown.
@@ III.2 Critical strip bath effects
We find significant differences in device performance and surface chemical content when using the two different resist strip baths for the fabrication. Figure 2(c) shows the results from dies using the 1165 strip bath with varying durations of post-fabrication acid treatment, and Fig. 3 shows the results from dies using the 1165 strip bath with no post-fabrication acid treatment, and dies using the AZ strip bath without and with acid treatment. For the recipe using 1165, a post-fabrication HF dip was necessary to achieve near-state-of-the-art (SOTA) performance. Without the final HF dip, the low and high power losses were found to be about 15 times higher with median low power losses: , , and for the BOE, Anneal, and Thermal samples, respectively (compare the 1165 samples from Fig. 3 to the acid-treated samples in Fig. 2(c)). At high power, the loss reduces only modestly (less than a factor of two) for the 1165 samples with no acid treatment, indicating the dominant change is a source of power-independent loss. Per Table 1, the Cl residue on the silicon surface (substrate-air interface) that results from the etch process is largely untouched by the strip bath but is removed by HF, as seen by XPS. However, the AZ bath successfully removes chlorine and other residues as well as substantially reducing surface carbon content. This correlates with performance equaling to the levels seen with 1165 with the following HF dip, as shown in Fig. 3. We tested with XPS two other NMP-based strip baths, and also found residual Cl on the Si surface (see Table 1) as well as elevated surface carbon content relative to AZ.
@@
The thicker surface oxide with the AZ300T samples motivated a return to attempting HF treatment following fabrication. With our usual treatment with 2% HF, the losses lowered further to a median value across three chips of DP2 devices, as shown in Fig. 3. These three samples were consistently below the 25% / 75% loss window of samples treated with HF after the 1165 strip, or with AZ300T alone. Moreover, these samples were all of the “DP2” chamber preparation, meaning that the Nb sputter target was co-loaded with magnetic targets that were used in depositions prior to the Nb depositions.
@@ IV Discussion
We also tested three different methods of preparing substrates, which did not yield clearly distinct measurements of device quality. This outcome is consistent with the metal-substrate interface being subdominant sources of low-power [woods_determining_2019, wenner_surface_2011] or the Nb-Si interface being largely unaffected by these differences in preparation. In order to further improve the quality of our devices we will consider other changes to our fabrication methods and materials stack synthesis of bilayer and alloy films to mitigate Nb oxide formation [bal_systematic_2023]. In the future, we will also explore co-loading superconductor and magnet sputter targets to investigate heterostructures for, e.g., -junctions [birge_ferromagnetic_2024, kim_superconducting_2024].
@@ Acknowledgements.
| Chamber | Prep | Strip | HF | Nc | Np | med | iqr | std |
| DP1 | BOE | 1165 | N | 1 | 7 | 31.1 | 10.1 | 8.0 |
| DP1 | Anneal | 1165 | N | 1 | 8 | 31.1 | 7.5 | 16.4 |
| DP1 | Thermal | 1165 | N | 1 | 6 | 33.9 | 11.6 | 28.9 |
| DP1 | BOE | 1165 | Y | 4 | 29 | 1.56 | 2.36 | 2.53 |
| DP1 | Anneal | 1165 | Y | 5 | 34 | 1.53 | 1.10 | 2.55 |
| DP1 | Thermal | 1165 | Y | 5 | 38 | 1.21 | 0.61 | 0.85 |
| DP1 | BOE | AZ | N | 2 | 11 | 1.08 | 0.33 | 0.25 |
| DP2 | BOE | AZ | N | 3 | 19 | 1.50 | 0.40 | 0.94 |
| DP2 | BOE | AZ | Y | 3 | 22 | 0.79 | 0.19 | 0.19 |
| LPS | BOE | AZ | N | 1 | 8 | 1.06 | 0.17 | 0.17 |
| LPS | Anneal | AZ | N | 1 | 8 | 1.18 | 0.21 | 0.93 |
@@ A.1 Chamber preparation
For deposition procedure 2 (DP2) we deposited Nb films while co-loaded targets included permalloy (Ni80Fe20), cobalt (Co), and cobalt-iron-boron (Co20Fe60B20) targets.
@@ A.2 Substrate preparation
| c (M-B) | |||
| (Median) | 1.48 | 1.53 | 1.21 |
The BOE preparation begins with an RCA clean with SC-1 and SC-2. The RCA Clean is a two-step clean: SC-1 and SC-2. SC-1 is a 10-min soak in a 6:1:1 mixture of DI water, ammonium hydroxide, and hydrogen peroxide, respectively. The clean is done at C, followed by a 10-min rinse in DI water. SC-2 is a 10-min soak at C in a 6:1:1 mixture of DI water, hydrochloric acid, and hydrogen peroxide, respectively. The clean is followed a 10-min water rinse and nitrogen gas dry. Within one hour fo completing the RCA clean the silicon oxide is removed by a 10:1 buffered oxide etch (BOE) soak for 1 minute, two 30-second deionized water rinses, and a blow dry with nitrogen. The wafer is loaded into the deposition chamber load lock within 5 min of concluding the BOE etch process. X-ray photoelectron spectroscopy (XPS) shows no significant oxide regrowth within 1 hour of the BOE etch, as shown in S5(b). We pump on the load lock for 1.5-2 hours, reaching pressures below Torr before transferring the wafer into the chamber for deposition.
@@
The Thermal preparation includes removal of the top layer of material from the wafer by a method adapted from [jakob_influence_1991]. For the thermal oxide growth, we first “cleaned” the furnace by running a 30 min, dry oxidation with an additional flow of hydrochloric gas on an empty chamber. The purpose of this is to remove ionic compound residue from the chamber. After this, the chamber was loaded with substrates, which were just cleaned with the two-step RCA clean. The dry thermal oxidation (without the use of HCl) was run for 6 hours resulting in about of silicon oxide [mcguire_semiconductor_1988]. The oxidation was followed by a 30 min nitrogen anneal at , done in the same chamber. The annealing step helps to create a uniform interface between the substrate and thermal silicon oxide. We begin with an RCA clean, followed by (dry) thermal oxidation of the wafers at for 6 hours, resulting in about of silicon oxide [mcguire_semiconductor_1988]. This layer is then removed by a 10:1 BOE soak, resulting in the removal of the top of silicon oxide, corresponding to roughly of the un-oxidized Si. After this, the wafer is prepared with the Anneal procedure.
@@ A.4 Niobium e-beam deposition
The e-beam deposited films were completed from a 99.95% pure Nb starter source loaded into a 40cc UHV single pocket electron beam source. The growth chamber base pressure is less than and the growth rate was . Prior to growth the 3-inch-diameter, float-zone-refined Si(001) wafer with resistivity was degreased with sequential baths of acetone, methanol, and isopropyl alcohol, and the native oxide was removed by a 5% hydroflouric acid etch for , followed by a deionized water rinse. The wafer was loaded into the HV load lock and heated to for to desorb water before transferring into the growth chamber. Niobium was deposited onto the substrate at . The first film was grown on the wafer as transferred. The second wafer was heated in-situ to for using a ramp rate of and then cooled to for growth. Reflection high energy electron diffraction was completed after growth showing rings of a polycrystalline film for the sample non-annealed substrate, but spotty/streaky pattern for the annealed sample. The in situ annealed sample also expressed weak Nb(110) peaks in the symmetric x-ray diffraction scan, while the sample that was not annealed did not express any Nb diffraction peaks. Atomic force microscopy measured an rms roughness of 1.2 nm for the non-annealed sample and 0.96 nm for the annealed sample using scans.
@@ C.1 Sample preparation
The HF dip of the 1165 and the AZ300T samples took place one month after the initial measurement. To ensure consistency in measurement, the sample was remeasured prior to the HF dip (1165 + 30D & AZ300T + 30D). The sample was then dipped in 2%HF for 60s and measured (1165 + HF & AZ300T + HF).
@@ D.3 Temperature dependence
where, is the effective superconducting gap, is the overall amplitude proportional to the kinetic inductance ratio, is the Boltzmann constant, and is the zero-th order modified Bessel function of the second kind. The average for each set of samples is reported in Table S2. We find that the of the three sample Nb resonators plotted in Figure S10 are , , and , for BOE, Anneal, and Thermal, respectively. The M-B values vary somewhat from the low-frequency measurements, which can be explained by the incomplete fit for the temperature dependent data.

### v2

@@ Abstract
Magnetic impurities are known to degrade superconductivity. For this reason, physical vapor deposition chambers that have previously been used for magnetic materials have generally been avoided for making high-quality superconducting resonator devices. In this article, we show by example that such chambers can be used for this purpose; with Nb films sputtered in a chamber that continues to be used for magnetic materials, we demonstrate compact ( gap) coplanar waveguide resonators with low-power internal quality factors near one million. We achieve this using a resist strip bath with no post-fabrication acid treatment, which results in performance comparable to previous strip baths with acid treatments. We also find evidence that this improved resist strip bath provides a better surface chemical template for post-fabrication hydrogen fluoride processing. These results are consistent across three Si substrate preparation methods, including a anneal. These results will inform nanofabrication for other superconducting materials and the integration of magnetic materials for hybrid systems.
@@ I Introduction
Superconducting microwave circuits are one of the leading candidate platforms for enabling quantum computing technology [kjaergaard_superconducting_2020, sivak_real-time_2023, acharya_quantum_2024], and Nb on Si is a commonly implemented materials set [mcrae_materials_2020]. A promising avenue for improving device performance and technological prospects is to eliminate defects from the materials and interfaces within the superconducting circuits [mcrae_materials_2020, de_leon_materials_2021].
@@ III.1 Si preparation series
We use the open-source resonator design developed by the Boulder Cryogenic Quantum Testbed to characterize the microwave properties of resonators fabricated from these films [kopas_simple_2022]. This design has a gap between the center conductor and the ground plane and has a hangar resonator configuration, with eight resonators at frequencies between 4 and . For each resonance, we locally vary the probe frequency and fit the complex-valued transfer function to the expected resonance response, as in Figure 2(b). Accomplishing this as a function of drive power, and using the measured attenuation of the input coaxial lines, we can estimate the average photon number. We fit the value of the internal quality factor (details in Appendix D.2) and calculate the loss ():
@@
In Figure 2(c), we show the internal loss of a resonator as a function of the average photon number in the resonator for devices corresponding to our three substrate preparation methods. The three sample resonators plotted are representative of the qualitatively different trends observed. In particular, the data from resonator 2 and resonator 3 show signs of a ‘double-S’ shaped curve, possibly indicative of two TLS (two level system) baths existing at different critical photon numbers [verjauw_investigation_2021]. This suggests that a usual TLS-based model is unsuitable for our data. Note that the data from resonator 1 does not have a second plateau present in the 1 to 100 photon range. Furthermore, none of the resonators loss curves saturate at high powers. One possible explanation is the renormalization of the critical exponent for the TLS bath due to TLS-TLS interactions, but this does not completely avoid a saturation effect [crowley_disentangling_2023, altoe_localization_2022, verjauw_investigation_2021]. Another possibility is a large number of TLS baths with a wide distribution of critical photon numbers. Given these discrepancies with usual models, we do not attempt to fit to those models.
@@
The results of the substrate preparation series are summarized in Figure 2(d). We believe that the distributions across all the dies from the different substrate preparation methods are not sufficiently well-separated to draw a clear distinction. As Nb resonators are susceptible to significant losses from the reformation of Nb oxide [zheng_nitrogen_2022, verjauw_investigation_2021, altoe_localization_2022, woods_determining_2019], some of the variation in the data may be explained by the varying amount of Nb re-oxidation after the HF treatment. In addition, temporal fluctuations of measured quality factor during the same cooldown for Nb on Si were found to have a relative standard deviation of 13% [vallieres_loss_2024], much smaller than what we observe cooldown-to-cooldown.
@@ III.2 Critical strip bath effects
We find significant differences in device performance and surface chemical content when using the two different resist strip baths for the fabrication. Figure 2(d) shows the results from dies using the 1165 strip bath with varying durations of post-fabrication acid treatment, and Fig. 3 shows the results from dies using the 1165 strip bath with no post-fabrication acid treatment, and dies using the AZ strip bath without and with acid treatment. For the recipe using 1165, a post-fabrication HF dip was necessary to achieve near-state-of-the-art (SOTA) performance. Without the final HF dip, the low and high power losses were found to be about 15 times higher with median low power losses: , , and for the BOE, Anneal, and Thermal samples, respectively (compare the 1165 samples from Fig. 3 to the acid-treated samples in Fig. 2(d)). At high power, the loss reduces only modestly (less than a factor of two) for the 1165 samples with no acid treatment, indicating the dominant change is a source of power-independent loss. Per Table 1, the Cl residue on the silicon surface (substrate-air interface) that results from the etch process is largely untouched by the strip bath but is removed by HF, as seen by XPS. However, the AZ bath successfully removes chlorine and other residues as well as substantially reducing surface carbon content. This correlates with performance equaling to the levels seen with 1165 with the following HF dip, as shown in Fig. 3. We tested with XPS two other NMP-based strip baths, and also found residual Cl on the Si surface (see Table 1) as well as elevated surface carbon content relative to AZ.
@@
The thicker surface oxide with the AZ300T samples motivated a return to attempting HF treatment following fabrication. With our usual treatment with 2% HF, the losses lowered further to a median value across three chips, as shown in Fig. 3. These three samples were consistently below the 25% / 75% loss window of samples treated with HF after the 1165 strip, or with AZ300T alone.
We also examined samples with a 10:1 post-BOE treatment, as an alternative to the post-HF treatment. We found a median value of for , as shown in Fig. 3. This performance exceeded that of the post-HF treatment, which could be indicative of hydrogen contamination in the HF-treated samples as previously reported [altoe_localization_2020, torres-castanedo_formation_2024]. In addition, we re-measured these two samples after 4 and 6 weeks of exposure to ambient conditions without removal from the packages. The aging process allows for both the SiO2 and NbOx to regrow [verjauw_investigation_2021]. We found that median after aging , returning to the range around the median value of for samples without any post-acid treatment (see Fig. 3).
@@ IV Discussion
We also tested three different methods of preparing substrates, which did not yield clearly distinct measurements of device quality. This outcome is consistent with the metal-substrate interface being subdominant sources of low-power loss [woods_determining_2019, wenner_surface_2011] or the Nb-Si interface being largely unaffected by these differences in preparation. In order to further improve the quality of our devices we will consider other changes to our fabrication methods and materials stack synthesis of bilayer and alloy films to mitigate Nb oxide formation [bal_systematic_2023]. In the future, we will also explore co-loading superconductor and magnet sputter targets to investigate heterostructures for, e.g., -junctions [birge_ferromagnetic_2024, kim_superconducting_2024].
@@ Acknowledgements.
| Chamber | Prep | Strip | Acid | Nc | Np | med | iqr | std |
| DP1 | BOE | 1165 | N | 1 | 7 | 29.7 | 7.4 | 4.2 |
| DP1 | Anneal | 1165 | N | 1 | 8 | 31.1 | 7.2 | 11.3 |
| DP1 | Thermal | 1165 | N | 1 | 6 | 32.1 | 7.1 | 19.6 |
| DP1 | BOE | 1165 | HF | 4 | 29 | 1.56 | 2.35 | 2.53 |
| DP1 | Anneal | 1165 | HF | 5 | 31 | 1.57 | 0.93 | 2.53 |
| DP1 | Thermal | 1165 | HF | 5 | 33 | 1.20 | 0.52 | 0.70 |
| DP1 | BOE | AZ | N | 5 | 30 | 1.45 | 0.65 | 0.47 |
| DP2 | BOE | AZ | N | 2 | 13 | 1.76 | 0.32 | 1.09 |
| DP1 | BOE | AZ | HF | 2 | 14 | 0.841 | 0.198 | 0.264 |
| DP2 | BOE | AZ | HF | 1 | 7 | 0.762 | 0.158 | 0.116 |
| DP1 | BOE | AZ | BOE | 1 | 7 | 0.501 | 0.150 | 0.262 |
| DP2 | BOE | AZ | BOE | 1 | 6 | 0.476 | 0.046 | 0.049 |
| DP1 | BOE | AZ | Aged | 1 | 5 | 1.64 | 0.71 | 0.36 |
| DP2 | BOE | AZ | Aged | 1 | 2 | 1.18 | 0.08 | 0.08 |
| LPS | BOE | AZ | N | 1 | 8 | 1.06 | 0.27 | 0.19 |
| LPS | Anneal | AZ | N | 1 | 8 | 1.20 | 0.28 | 0.90 |
@@ A.1 Chamber preparation
For deposition procedure 2 (DP2) we deposited Nb films while co-loaded targets included permalloy (Ni80Fe20), cobalt (Co), and cobalt-iron-boron (Co20Fe60B20).
@@ A.2 Substrate preparation
| Tc (M-B) | |||
| (Median) | 1.56 | 1.57 | 1.20 |
The BOE preparation begins a two-step RCA Clean: SC-1 and SC-2. SC-1 is a 10-min soak in a 6:1:1 mixture of DI water, ammonium hydroxide, and hydrogen peroxide, respectively. The clean is done at C, followed by a 10-min rinse in DI water. SC-2 is a 10-min soak at C in a 6:1:1 mixture of DI water, hydrochloric acid, and hydrogen peroxide, respectively. The clean is followed a 10-min water rinse and nitrogen gas dry. Within one hour of completing the RCA clean the silicon oxide is removed by a 10:1 buffered oxide etch (BOE) soak for 1 minute, two 30-second deionized water rinses, and a blow dry with nitrogen. The wafer is loaded into the deposition chamber load lock within 5 min of concluding the BOE etch process. X-ray photoelectron spectroscopy (XPS) shows no significant oxide regrowth within 1 hour of the BOE etch, as shown in S5(b). We pump on the load lock for 1.5-2 hours, reaching pressures below Torr before transferring the wafer into the chamber for deposition.
@@
The Thermal preparation includes removal of the top layer of material from the wafer by a method adapted from [jakob_influence_1991]. For the thermal oxide growth, we first “cleaned” the furnace by running a 30 min, dry oxidation with an additional flow of hydrochloric gas on an empty chamber. The purpose of this is to remove ionic compound residue from the chamber. After this, the chamber was loaded with substrates, which had just cleaned with the two-step RCA clean. The dry thermal oxidation (without the use of HCl) was run for 6 hours resulting in about of silicon oxide [mcguire_semiconductor_1988]. The oxidation was followed by a 30 min nitrogen anneal at , done in the same chamber. The annealing step helps to create a uniform interface between the substrate and thermal silicon oxide. This layer is then removed by a 10:1 BOE soak, resulting in the removal of the top of silicon oxide, corresponding to roughly of the un-oxidized Si. After this, the wafer undergoes the flash anneal and cooldown described as part of the Anneal preparation.
@@ A.4 Niobium e-beam deposition
The e-beam deposited films were completed from a 99.95% pure Nb starter source loaded into a 40cc UHV single pocket electron beam source. The growth chamber base pressure is less than and the growth rate was . Prior to growth the 3-inch-diameter, float-zone-refined Si(001) wafer with resistivity was degreased with sequential baths of acetone, methanol, and isopropyl alcohol, and the native oxide was removed by a 5% hydroflouric acid etch for , followed by a deionized water rinse. The wafer was loaded into the HV load lock and heated to for to desorb water before transferring into the growth chamber. Niobium was deposited onto the substrate at . The first film was grown on the wafer as transferred. The second wafer was heated in-situ to for using a ramp rate of and then cooled to for growth. Reflection high energy electron diffraction was completed after growth showing rings of a polycrystalline film for the sample non-annealed substrate’s sample, but a spotty/streaky pattern for the annealed substrate’s sample. Also, the sample with in situ annealed substrate expressed weak Nb(110) peaks in the symmetric x-ray diffraction scan, while the sample without the annealed substrate did not express any Nb diffraction peaks. Atomic force microscopy measured an rms roughness of for the Nb film with non-annealed substrate and for the annealed substrate using × scans.
@@ C.1 Sample preparation
The HF dip of the 1165 and the AZ300T samples took place one month after the initial measurement. To ensure consistency in measurement, the sample was remeasured prior to the HF dip (1165 + 41D & AZ300T + 51D). The sample was then dipped in 2%HF for 60s and measured (1165 + HF & AZ300T + HF).
@@ D.3 Temperature dependence
where, is the effective superconducting gap for elemental Nb, is the overall amplitude proportional to the kinetic inductance ratio, is the Boltzmann constant, and is the zero-th order modified Bessel function of the second kind. The average for each set of samples is reported in Table S2. We find that the of the three sample Nb resonators plotted in Figure S10 are , , and , for BOE, Anneal, and Thermal, respectively. The M-B values vary somewhat from the low-frequency measurements, which can be explained by the incomplete fit for the temperature dependent data.